For a particular RLC series circuit, the capacitive reactance is 6.85 Ω, the inductive reactance is 36.7 Ω, and the maximum voltage across the 92.5-Ω resistor is 35.1 V. What is the maximum voltage across the circuit?
Answer #1

Solution:

Given -

XC = 6.85 ohm

XL = 36.7 ohm

R = 92.5 ohm

VR = 35.1 V

IR = VR / R = 35.1 / 92.5 = 0.379 A

Here, Impedance: Z = sqrt { R2 + ( XL - XC)2} = 97.197 ohm

the maximum voltage across the circuit , V = IR * Z = (0.379 A) x (97.197 Ohm) = 36.88 V

Therefore: the maximum voltage across the circuit is 36.88 V